It takes 9 hours to fill a pool with the inlet pipe. It can be emptied in 12 hours with the outlet pipe. If the pool is 1/2 full to begin with, how long will it take to fill it from there if both pipes are open?
Answer by richwmiller(17219)

The student says 1/2 full.
1/9-1/12=1/2x
12/108-9/108=1/2x
3/108=1/2x
2/36=1/x
1/18=1/x
x=18
If it were empty
1/9-1/12=1/x
3/108=1/x
x=36